OpenStudy (anonymous):

solve the equation 6x - 8y = 72

OpenStudy (anonymous):

in terms of which variable?

OpenStudy (saifoo.khan):

u want to graph?

OpenStudy (anonymous):

the actual question is select a solution to the equation 6x - 8y = 72 A. (8,1) B. (8, -1) C. (8,-3) OR D. (8,3)

OpenStudy (anonymous):

substitute the values 6(8)-8(-3) = 72 48+24=72 72=72 choice C

OpenStudy (saifoo.khan):

Yes, he's correct.